# Turbulence to Order Transitions in Activity Patterned Active Nematics

### v1

@@ Abstract
To further probe the turbulent to vortex transition, we simulate a system in which the activity is confined to circles that are arranged in a triangular lattice, as shown in Fig. 1(b). The area fraction of the active regions is given by $\nu=(\pi/2\sqrt{3})(2R/a)^{2}$, where $R$ is the circle radius and $a$ is the lattice constant. We vary $a/2R$ to give values of $\nu$ in the range $\nu=0.1$ to $\nu=0.7$. Even without the 1D stripe arrangement of activity, we observe a transition from active turbulence to a stable vortex state when $\alpha_{0}$ and $a/2R$ are increased. In Fig. 3(a) we plot a time snapshot of the vorticity and velocity fields for a simulation with $a/2R=1.28$ and $\alpha_{0}=0.75$. Here we observe active turbulence in which the locations of the vortices are not correlated with the locations of the activity spots. When we increase $\alpha_{0}$ and $a/2R$, as shown in Fig. 3(b) for $\alpha_{0}=1.25$ and $a/2R=1.74$, individual vortices become localized in each active region. In Fig. 3(c), we plot a heat map of $\log|K|$ measured over the entire sample as a function of $\alpha_{0}$ versus $a/2R$. Below the line marking points at which $\log|K|=0$ the system exhibits active turbulence, and above the line the system relaxes to a configuration with a stable vortex at each active site. The turbulent to vortex transition has a similar form to what we observed in the active stripe sample from Fig. 2(e,f), but the 1D active turbulent state is lost due to the geometry of the active circles.
@@
It is surprising that for both the active stripes and the active circles, the transition to the turbulent state shifts to smaller spacings between the active regions (and a greater total active area) when the activity level $\alpha_{0}$ is increased since, in confined systems with homogeneous activity, an increase in the activity typically produces an increased amount of topological defect unbinding and a more rapid transition to turbulence [4, 51, 6]. In our systems with inhomogeneous activity, we can define an active length scale $\ell_{\alpha}\propto\sqrt{1/\alpha_{0}}$ and a characteristic length scale $\ell_{s}$ of the active regions. We find that when $\ell_{\alpha}\gtrsim\ell_{s}$, the system exhibits active turbulence, while when $\ell_{\alpha}\lesssim\ell_{s}$, localized vortical structures appear [45]. This has implications even for systems that do not have explicitly patterned activity, since in any experimental system the activity is not distributed homogeneously. Our results indicate that the approximation of homogeneously distributed activity is valid in the limit of small $\ell_{s}$ and small $\alpha_{0}$, but outside this limit, the effect of inhomogeneous activity may disrupt the active turbulent phase.
A more careful inspection of the vortex state that appears in the active stripe system for large $\alpha_{0}$, as in Fig. 2(d), reveals antiferromagnetic ordering of the vortices along the $y$ direction and ferromagnetic ordering along the $x$ direction. To quantify this ordering, we compute the autocorrelation functions of the vorticity along the $x$ and $y$ directions at various values of $y_{0}$ and $x_{0}$:
$C_{y}(x_{0},Y,t)$ $=\frac{\langle\omega(x_{0},y+Y,t)\omega(x_{0},y,t)\rangle_{y}}{\langle\omega(x_{0},y,t)^{2}\rangle_{y}}.$ (6)
We define $\chi_{x}(t)=\langle C_{x}(\ell_{s}+\ell_{i},y_{0},t)\rangle_{y_{0}}$ and $\chi_{y}(t)=\langle C_{y}(x_{0},\ell_{h},t)\rangle_{x_{0}}$ so that $\chi_{x}$ ($\chi_{y}$) measures the nearest neighbor correlation in the $x$ ($y$) direction. Here, $\chi_{x,y}\in[-1,1]$, with positive (negative) values indicating ferromagnetic (antiferromagnetic) order and larger magnitudes indicating stronger order. An example of the calculation of $\chi_{x}$ and $\chi_{y}$ appears in the Supplemental Material [45].
@@
We compute $\chi_{x,y}$ for $\alpha_{0}>1.2$, where $\log|K|$ is large and positive. In Fig. 4(a) at $\ell_{i}=20$ and $\alpha_{0}=1.5$, for long times $\chi_{x}\approx+0.5$ and $\chi_{y}\approx-1$, indicating moderate ferromagnetic order along the $x$ direction and strong antiferromagnetic order along the $y$ direction. For smaller $\ell_{i}$, $\chi_{x}$ and $\chi_{y}$ fluctuate more over time, as shown in Fig. 4(b) at $\ell_{i}=8$ and $\alpha_{0}=1.4$. Here, $\chi_{x}$ fluctuates around zero but $\chi_{y}$ fluctuates around $\chi_{y}=-0.5$, indicating a lack of ordering along the $x$ direction and moderate antiferromagnetic order along the $y$ direction. In Figs. 4(c,d) we plot heat maps of the time averaged $\langle\chi_{x}\rangle$ and $\langle\chi_{y}\rangle$, respectively, as a function of $\alpha_{0}$ versus $\ell_{i}$. As $\ell_{i}$ increases, $\langle\chi_{x}\rangle$ grows from roughly zero to $\approx 0.6$, indicating the emergence of ferromagnetic order transverse to the activity stripes. In contrast, $\langle\chi_{y}\rangle$ is almost always negative, and reaches values very close to $-1$ when $\ell_{i}>14$, indicating almost perfect, stable antiferromagnetic order along the active stripe.
Antiferromagnetic vortex ordering has been observed for active nematics confined to channels, as well as in two-dimensional systems with increasing substrate friction or obstacle confinement [51, 52, 22, 25, 32, 36]. The unbinding of topological defects helps promote antiferromagnetic order when vortices are stable, since the negative defects can sit between the vortices and mediate a rotation of the nematic director as the chirality of the positive defects at the center of each vortex alternates. A time average of the topological charge density in our active stripe system reveals negative defects sitting between the vortices for simulations in the large $\ell_{i}$ and large $\alpha_{0}$ regime [45]. The ferromagnetic order transverse to the active regions is of great interest since, to our knowledge, it has only been reported numerically for active nematics interacting with periodic arrays of obstacles [32], while vortex synchronization has only recently been reported in acoustically actuated active nematics [12]. For well separated active stripes, the only interaction that couples the stripes is the elastic interaction of the nematic, which has been shown to promote ferromagnetic order of vortices in obstacle laden systems [32]. We therefore conclude that the ferromagnetic order transverse to the stripes is driven by elastic interactions, while the antiferromagnetic order along the stripes is driven by active forces.
@@
Figure 5: (a) Time snapshot of $\omega/|\omega_{max}|$ (color) and $\langle{\bf v}\rangle$ (arrows) for lines of activity circles (bright spots) with $\ell_{i}=14$ and $\alpha_{0}=1.5$. (b) The corresponding time-averaged topological defect density. (c) $\langle\chi_{x}\rangle$ and (d) $\langle\chi_{y}\rangle$ vs $\ell_{i}$ for systems of activity stripes (blue) and circle lines (red) with $\alpha_{0}=1.5$.
@@
Finally, we investigate the impact of the 1D degree of freedom along the stripes on the vortex ordering. While the antiferromagnetic order develops quickly along the active stripes, the ferromagnetic order takes much longer to develop. Qualitatively, the vortices slide into place over time in order to develop ferromagnetic order along the $x$ direction, rather than nucleating in ordered positions initially. To remove the 1D degree of freedom, we simulate lines of active circles [Fig. 1(c)] for $\alpha_{0}=1.5$ at varied $\ell_{i}$. As shown in Fig. 5(a), vortices quickly develop in the active regions, while Fig. 5(b) indicates that positive topological defects rotate in the center of the vortices while negative defects sit between active regions. Although this is similar to the behavior of the active stripe system, the vortex ordering is different. In Figs. 5(c) and (d) we plot $\langle\chi_{x}\rangle$ and $\langle\chi_{y}\rangle$ versus $\ell_{i}$ for the stripe and circle line systems. The measures diverge for large $\ell_{i}$, with $\langle\chi_{x}\rangle\approx 0$ and $\langle\chi_{y}\rangle\gtrsim 0$ for the circle lines, but both measures reaching nonzero values for the stripe system. The vortices develop quickly due to the large active forces in the active regions, but in the circle line system, they remain frozen in place and cannot reorder along $x$ [45]. Ordinarily the vortices would naturally adopt an antiferromagnetic ordering along $y$ in a 1D channel, as in the stripe system, but this ordering requires an odd number of negative topological defects to sit between pairs of vortices to mediate the change in chirality. As shown in Fig. 5(b), in the circle line system, the divots between adjacent circles permit two defects to sit between vortices along the edge of the inactive region. We find that when one defect is present, the chirality of neighboring vortices changes, but when two defects are present, the chirality remains the same. This result underscores the pivotal role that activity geometry plays in determining the ordering of the system.
@@
Conclusion– We have studied the effects of periodic, inhomogeneous activity patterning on active nematics. We observe two non-equilibrium transitions for a system of activity stripes: a transition from 2D to 1D active turbulence, and a transition from 1D active turbulence to a well-defined vortical flow state that can exhibit ferromagnetic and antiferromagnetic vortex ordering. The vortex ordering can be destroyed if the activity stripes are replaced by lines of activity circles. When we introduce a triangular array of activity circles, the transition to a turbulent state is suppressed if the activity is too large due to the tendency of localized vortices to emerge at the activity spots. Our results provide a method for inducing and controlling non-equilibrium phase transitions in active nematics and for producing customized activity patterned ordered states.

### v2

@@ III.2 Activity Circles
To further probe the turbulent to vortex transition, we simulate a system in which the activity is confined to circles that are arranged in a triangular lattice, as shown in Fig. 1(b). The area fraction of the active regions is given by $\nu=(\pi/2\sqrt{3})(2R/a)^{2}$, where $R$ is the circle radius and $a$ is the lattice constant. We vary $a/2R$ to give values of $\nu$ in the range $\nu=0.1$ to $\nu=0.7$. Even without the 1D stripe arrangement of activity, we observe a transition from active turbulence to a stable vortex state when $\alpha_{0}$ and $a/2R$ are increased. In Fig. 4(a) we plot a time snapshot of the vorticity and velocity fields for a simulation with $a/2R=1.28$ and $\alpha_{0}=0.75$. Here we observe active turbulence in which the locations of the vortices are not correlated with the locations of the activity spots. When we increase $\alpha_{0}$ and $a/2R$, as shown in Fig. 4(b) for $\alpha_{0}=1.25$ and $a/2R=1.74$, individual vortices become localized in each active region. In Fig. 4(c), we plot a heat map of $\log|K|$ measured over the entire domain as a function of $\alpha_{0}$ versus $a/2R$. Below the line marking points at which $\log|K|=0$ the system exhibits active turbulence, and above the line the system relaxes to a configuration with a stable vortex at each active site. The turbulent to vortex transition has a similar form to what we observed in the active stripe sample from Fig. 2(e,f), but the 1D active turbulent state is lost due to the geometry of the active circles.
@@
It is surprising that, for both the active stripes and the active circles, the transition to the turbulent state shifts to smaller spacings between the active regions (and a greater active area fraction) when the activity level $\alpha_{0}$ is increased since, in confined systems with homogeneous activity, an increase in the activity typically produces an increased amount of topological defect unbinding and a more rapid transition to turbulence [4, 55, 6]. In the systems with inhomogeneous activity studied here, we find that when $\ell_{\alpha}\gtrsim\ell_{i}$, the system exhibits active turbulence, while when $\ell_{\alpha}\lesssim\ell_{i}$, localized vortical structures appear. We show this in Fig. 5 by plotting the 2D active turbulence transition curves from both the active stripe and circle systems in Figs. 3(b) and 4(c) against axes $1/\sqrt{\alpha_{0}}$ and $\ell_{i}/\ell_{s}$. In the figure, we have identified $\ell_{i}/\ell_{s}=(a-2R)/2R$ for the activity circles. For both systems, the transitions depend linearly on $1/\sqrt{\alpha_{0}}\propto\ell_{\alpha}$ as functions of $\ell_{i}/\ell_{s}$, indicating that the transitions are given by a direct interplay between the two length scales.
The dependence of the transition on activity density has implications even for systems that do not have explicitly patterned activity, since in any experimental system the active forces will not be distributed homogeneously. Our results indicate that the approximation of homogeneously distributed activity is valid in the limit of $\ell_{\alpha}>>\ell_{i}$, but outside this limit, the effect of inhomogeneous activity may disrupt the active turbulent phase.
Figure 5: 2D active turbulence transition curves for activity stripes and activity circle lattice plotted against $1/\sqrt{\alpha_{0}}$ versus $\ell_{i}/\ell_{s}$. For the lattice of activity circles we have made the identification $\ell_{i}/\ell_{s}=(a-2R)/2R$.
IV Ordering of Vortices
IV.1 Activity Stripes
We now investigate the high activity region of the phase diagram of activity stripes, characterized by well-defined, slowly evolving vortices. A more careful inspection of this vortical flow state reveals strong antiferromagnetic ordering of the vortices along the $y$ direction and a tendency for ferromagnetic ordering along the $x$ direction. Using the one-dimensional correlation functions defined in Eqs. (16) and (17), $C_{x}(X,y_{0},t)$ and $C_{y}(x_{0},Y,t)$, we may quantify this ordering by defining $\chi_{x}(t)=\langle C_{x}(\ell_{s}+\ell_{i},y_{0},t)\rangle_{y_{0}}$ and $\chi_{y}(t)=\langle C_{y}(x_{0},\ell_{h},t)\rangle_{x_{0}}$ so that $\chi_{x}$ ($\chi_{y}$) measures the nearest neighbor correlation in the $x$ ($y$) direction. Here, $\chi_{x,y}\in[-1,1]$, with positive (negative) values indicating ferromagnetic (antiferromagnetic) order and larger magnitudes indicating stronger order. We show an example of the calculation of $\chi_{x}$ and $\chi_{y}$ for a frame from a simulation with $\ell_{i}=16$ and $\alpha_{0}=1.5$ in Fig. 6. We select $y_{0}=0$ and $x_{0}=-1.5(\ell_{i}+\ell_{s})=-39$, as illustrated in Fig. 6(a). Figures 6(b,c) show the corresponding $C_{x,y}$ versus $X$ and $Y$. For $C_{x}$ the peaks of the correlation function naturally occur at multiples of $\ell_{s}+\ell_{i}$ since the centers of the activity regions are separated by this distance, while the peaks of $C_{y}$ naturally occur at multiples of $\ell_{h}$ since this is the characteristic size of vortices. Thus, the first peak for each of these functions measures the nearest neighbor correlations [note that $C_{y}$ in Fig. 6(c) alternates between negative and positive peaks because of the antiferromagnetic order]. We average over multiple values of $y_{0}$ and $x_{0}$ to get the values of $\chi_{x}$ and $\chi_{y}$ presented below.
Figure 6: Graphic example of the calculation of $\chi_{x}$ and $\chi_{y}$. (a) Frame of the vorticity field for a simulation with $\ell_{i}=16$ and $\alpha_{0}=1.5$. The orange lines indicate $x_{0}=-39$ and $y_{0}=0$. (b) Vorticity autocorrelation function along the $x$ direction $C_{x}(X,y_{0})$. The value of the first peak gives $\chi_{x}$. (c) $C_{y}(x_{0},Y)$. The value of the first peak gives $\chi_{y}$.
We compute $\chi_{x,y}$ for $\alpha_{0}>1.2$, where $\log|K|$ is large and positive. We note that below $\alpha_{0}<1.2$ the system is in either 1D or 2D active turbulence, and the vortices are not stable enough to yield meaningful measurements of $\chi_{x,y}$. We plot $\chi_{x}(t)$ and $\chi_{y}(t)$ for a system with $\ell_{i}=20$ and $\alpha_{0}=1.5$ in Fig. 7(a). For long times, after the system reaches a stationary state, we find $\chi_{x}\approx+0.5$ and $\chi_{y}\approx-1$, indicating moderate ferromagnetic order along the $x$ direction and strong antiferromagnetic order along the $y$ direction. For smaller $\ell_{i}$, $\chi_{x}$ and $\chi_{y}$ fluctuate more over time, as shown in Fig. 7(b) at $\ell_{i}=8$ and $\alpha_{0}=1.4$. Here, $\chi_{x}$ fluctuates around zero but $\chi_{y}$ fluctuates around $\chi_{y}=-0.5$, indicating a lack of ordering along the $x$ direction and moderate antiferromagnetic order along the $y$ direction. In Figs. 7(c,d) we plot heat maps of the time averaged $\langle\chi_{x}\rangle$ and $\langle\chi_{y}\rangle$, respectively, as a function of $\alpha_{0}$ versus $\ell_{i}$. As $\ell_{i}$ increases, $\langle\chi_{x}\rangle$ grows from roughly zero to $\approx 0.6$, indicating the emergence of ferromagnetic order transverse to the activity stripes. In contrast, $\langle\chi_{y}\rangle$ is almost always negative, and reaches values very close to $-1$ when $\ell_{i}>14$, indicating almost perfect, stable antiferromagnetic order along the activity stripes.
@@ IV.1 Activity Stripes
Antiferromagnetic vortex ordering has been observed for active nematics confined to channels, as well as in two-dimensional systems with increasing substrate friction or obstacle confinement [55, 58, 22, 25, 32]. The unbinding of topological defects helps promote antiferromagnetic order when vortices are stable, since the negative defects can sit between the vortices and mediate a rotation of the nematic director as the chirality of the positive defects at the center of each vortex alternates. We can see this in the time averaged plots of $D$ in Fig. 2(d) where the positive defects stay near the center of stable vortices, while single negative defects sit between them. We note that even though positive defects are typically motile in active nematics, they may form bound states of co-rotation at high activities [62, 63]. In the supplemental movies for simulations in this regime [56], the positive defects sometimes appear as defects with $+1$ winding at the center of the fluid vortices. This is likely due to the finite resolution of the simulation mesh, since we see in Fig 2(d) that the average defect density forms a ring around the center of the vorticies, rather than being concentrated at the center, which indicates that two defects are co-rotating in this region.
The ferromagnetic order transverse to the active regions is of great interest since, to our knowledge, it has only been reported numerically for active nematics interacting with periodic arrays of obstacles [32], while vortex synchronization has only recently been reported in acoustically actuated active nematics [12]. For well separated active stripes, the only interaction that couples the stripes is the elastic interaction of the nematic, which has been shown to promote ferromagnetic order of vortices in obstacle laden systems [32]. We therefore conclude that the ferromagnetic order transverse to the stripes is driven by elastic interactions, while the antiferromagnetic order along the stripes is driven by active forces.
@@
IV.2 Comparison to columns of active circles
@@
Finally, we investigate the impact of the 1D degree of freedom along the stripes on the vortex ordering. While the antiferromagnetic order develops quickly along the active stripes, the ferromagnetic order takes much longer to develop. Qualitatively, the vortices slide into place over time in order to develop ferromagnetic order along the $x$ direction, rather than nucleating in ordered positions initially. We thus suspect that the 1D degree of freedom along the $y$ direction afforded by the stripes is important for observing ordering along the $x$ direction. To test this, we remove the 1D degree of freedom by simulating columns of active circles [Fig. 1(c)] for $\alpha_{0}=1.5$ at varied $\ell_{i}$. As shown in Fig. 8(a), single vortices develop in the active regions, similar to both the stripe and circle lattice systems at this activity. We also show in Fig. 8(b) that positive topological defects rotate in the center of the vortices while negative defects sit between active regions.
Although the vortex stabilizing behavior in the active circle columns is similar to that of the active stripe system, the vortex ordering is different. In Figs. 8(c) and (d) we plot $\langle\chi_{x}\rangle$ and $\langle\chi_{y}\rangle$ versus $\ell_{i}$ for the stripe and circle column systems. The measures diverge for large $\ell_{i}$, with $\langle\chi_{x}\rangle\approx 0$ and $\langle\chi_{y}\rangle\gtrsim 0$ for the circle columns, but both measures reaching nonzero values for the stripe system. We note that, for the active circle columns, the averages $\langle\chi_{x,y}\rangle$ are averaged over time and three separate realizations with different random initial conditions. The vortices develop quickly due to the large active forces in the active regions, but in the circle column system, they remain frozen in place and do not reorder along $x$. An example of the dynamics is shown in Supplementary Movie 8.
The measures of $\chi_{x}$ in Fig. 8(c) confirm our hypothesis that the 1D degree of freedom is important for observing ferromagnetic ordering along the transverse direction to the stripes. Interestingly, even though the vortices would naturally adopt an antiferromagnetic ordering along $y$ in a 1D channel, as in the stripe system, we observe no ordering, or even a slight ferromagnetic ordering when the columns are well separated in the circle column system. As we noted for the activity stripes, antiferromagnetic ordering requires an odd number of negative topological defects to sit between pairs of vortices to mediate the change in chirality. As shown in Fig. 8(b), in the circle column system, the divots between adjacent circles permit two defects to sit between vortices along the edge of the inactive region. We find that when one defect is present, the chirality of neighboring vortices changes, but when two defects are present, the chirality remains the same [compare Figs. 8(a) and 8(b)]. This result underscores the pivotal role that activity geometry plays in determining the ordering of the system, as well as the interplay between vortex patterning and topological defect dynamics in active nematic systems.
Figure 8: (a) Time snapshot of $\omega/|\omega_{max}|$ (color) and $\langle{\bf v}\rangle$ (arrows) for columns of activity circles (bright spots) with $\ell_{i}=14$ and $\alpha_{0}=1.5$. (b) The corresponding time-averaged topological defect density. (c) $\langle\chi_{x}\rangle$ and (d) $\langle\chi_{y}\rangle$ vs $\ell_{i}$ for systems of activity stripes (blue) and circle columns (red) with $\alpha_{0}=1.5$.
V Conclusion
@@
We have studied the effects of periodic, inhomogeneous activity patterning on active nematics. We observe two non-equilibrium transitions for a system of activity stripes: a transition from 2D to 1D active turbulence, and a transition from 1D active turbulence to a well-defined vortical flow state that can exhibit ferromagnetic and antiferromagnetic vortex ordering. When we compare the activity stripes to a triangular array of activity circles, the turbulence transitions depend on the interplay between the active length scale and the density of active regions in similar manners, indicating an independence of activity patterning geometry on the transition to turbulence. On the other hand, we show that vortex ordering depends strongly on patterning geometry by comparing the activity stripes to columns of activity circles, which destroy the observed antiferromagentic and ferromagnetic vortex order. Our results provide a method for inducing and controlling non-equilibrium phase transitions in active nematics and for producing customized activity patterned ordered states.
It has been recently shown that activity patterning is an experimental possibility [37, 38, 39, 40], thus our results should be readily amenable to experimental verification. An interesting direction for future studies would be to understand the additional role of the hydrodynamic screening length, which we kept fixed in this study. It is possible that additional non-equilibrium phases may be observed when this parameter is varied and made commensurate with other system length scales. Additionally, an alternative approach to introduce activity heterogeneity would be to consider a spatially varying substrate friction, which was recently explored experimentally [26]. In the model studied here, this would cause both parameters in Eq. (11) to be spatially varying, and could lead to interesting interplays between the screening length and activity. Another direction would be to study the effect of temporally evolving activity fields and how the structure of turbulence or order is preserved or destroyed by motion of the activity fields. Recent work has shown that there is an optimal “surfing” speed for a single stripe of activity [34], and it would be interesting to study how these results change for multiple stripes or circles of activity. Finally, while our work gives some insight into the experimentally realistic case of inhomogeneous activity distributions, more insight may be gleaned by studying random activity patterning, as opposed to the periodic patterning considered here.